The amount of time Jesse works on his homework each day varies inversely with the amount of time he spends watching TV that day. If he spends 3 hours working on homework when he watches 0.5 hours of TV, how much time does he spend on homework when he watches 2.5 hours of TV.
Please help THANKS! :)

Answer by stanbon(48558)

The amount of time Jesse works on his homework each day varies inversely with the amount of time he spends watching TV that day. If he spends 3 hours working on homework when he watches 0.5 hours of TV, how much time does he spend on homework when he watches 2.5 hours of TV.
homework time = k/(TV time)
Need to find "k".
3=k/0.5
k=1.5
Equation is now homework time=1.5/(TV time)
If he watches 2.5hrs of TV, then
homework time= 1.5/2.5
homework time= 0.6 hrs = 36 minutes
Cheers,
Stan H.